OpenStudy (anonymous):

Find the least common denominator (LCD) for the fractions. A. 2 B. 10 C. 30 D. 60

OpenStudy (anonymous):

what are your fractions?

OpenStudy (anonymous):

The easiest way to find the LCD is to multiply the denominators. So 6*10=?

OpenStudy (anonymous):

\[6\times10=60\]

OpenStudy (anonymous):

yes so the LCD is 60.

zepdrix (zepdrix):

That is a fine and dandy denominator :) It is not the `least` of which though.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Can you help with three more though? :))

zepdrix (zepdrix):

Look for what numbers they have in common: \(\large\rm 6=3\cdot\color{orangered}{2}\) \(\large\rm 10=5\cdot\color{orangered}{2}\) So your smallest denominator that can be made from 6 and 10 is going to be the result of \(\large\rm 3\cdot5\cdot2\)

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

If you want me to

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

7 goes in to 28 four times what is one fourth of forty?

OpenStudy (anonymous):

in order to get the 28 to 7 we had to divide it by 4 so we need to do the same for the denominators. 40/4=?

OpenStudy (anonymous):

\[40\div4=10\]

OpenStudy (anonymous):

So a?

OpenStudy (anonymous):

yes. very good

OpenStudy (anonymous):

Thanks 2 more! :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

in order to get the 42 to 7 we had to divide it by 6 so we need to do the same for the numerators. 54/6=?

OpenStudy (anonymous):

\[54\div6=9\]

OpenStudy (anonymous):

yes so your answer is 9 or D

OpenStudy (anonymous):

Yay! Thanks so much! :)

OpenStudy (anonymous):

Next one!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to find the answer to this you need to find the LCD so 12*20=240. Now we need to change the numerators. So since we multiplied the 12 by 20 we need to do the same with the numerator of that fraction so 7*20=140, so our first fraction is 140/240. Now for the second fraction we multiplied the 20 by 12 so we need to do the same to the numerator so 11*12=132, so our second fraction is 132/240 and 132/240. Which one the numerators is higher? The one with the higher numerator takes up more space and therefore is the larger one.

OpenStudy (anonymous):

So we know that they are not equal so C is out

OpenStudy (anonymous):

so uh A?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

No not A because that fraction became 140/240 and that is larger then the other fraction, which is 132/240, and A shows that the second fraction is larger.

OpenStudy (anonymous):

Oh, so A is out.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the only option is B.

OpenStudy (anonymous):

yes. which is correct because it shows that the first fraction is larger than the second fraction.

OpenStudy (anonymous):

sorry if I confused you on that one

OpenStudy (anonymous):

Thanks now i have one last questions :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

No, no you didn't confuse anyone thanks for helping me!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Well A and B are out because, while the denominators do equal 48 when multiplied together, that is actually not the LCD, 24 is.

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Now we know that we multiplied the 6 by 4 to get to 24 so we need to multiply the 5 by 4 which equals 20 so our first fraction is 20/24. Since we multiplied 8 by 3 to get 24 we need to multiply the 3 by 3 which equals 9 so our second fraction is 9/24.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

very good :)

OpenStudy (anonymous):

Yay! Thanks for all the help!! :) You're the best! :))

OpenStudy (anonymous):

Got a 90% :) thanks!

OpenStudy (anonymous):

Lol you're welcome. I'm glad that I could help :)

OpenStudy (anonymous):

which one did you miss? If you don't mind me asking

OpenStudy (anonymous):

Let me check, hold on. :)

OpenStudy (anonymous):

This one was incorrect, the first one.

OpenStudy (anonymous):

It was actually 30 but at least i got a good grade. :)

OpenStudy (anonymous):

Ok I am sorry that I got it wrong. Good job on your grade :)

OpenStudy (anonymous):

No need to be sorry because if you wouldn't have helped me i would've gotten a bad grade, thanks again i really appreciate it! :D

OpenStudy (anonymous):

You're welcome. Have a nice day :)

OpenStudy (anonymous):

You too! :) Thanks again!